# Data-Driven Spectral Analysis Through Pseudo-Resolvent Koopman Operator in Dynamical Systems

### v1

@@ 6.1 Pendulum
In Figure 2, we approximate the Koopman operator $\mathbf{K}_{N}$ via EDMD and the resolvent operator $\mathbf{R}(z,\mathbf{K}_{N})$ using our method with dictionary sizes of 441, 600, and 806 basis functions, respectively, on another simulated dataset. The grid points $z$ are chosen along the circle $\Gamma_{1.01}$, a slight perturbation of the unit circle to avoid singularity when computing $\|\mathbf{R}(z,\mathbf{K}_{N})\|$. As the dictionary size $N$ increases, $1/\|\mathbf{R}(z,\mathbf{K}_{N})\|$ becomes smaller for a larger number of grid point $z$, thus leading to more points being detected as part of the approximated spectrum along the circle $\Gamma_{1.01}$ for a fixed threshold. This trend suggests a convergence behavior as $N$ grows.
In Figure 3, We also have tested for the generator as computed in Eq.(6). In continuous time case, since the spectrum of the Koopman operator is the unit circle, the corresponding spectrum for the Koopman generator $\mathcal{A}$ should be the imaginary axis. In the first figure, the spectrum is computed by gEDMD [KLUS2020132416]. Since our basis functions, i.e., the Fourier basis, are selected from the frequency close to 0, this explains that the evaluation for $z$ closer to 0 is more accurate, in other words, $1/\|\mathbf{R}(z,\mathbf{A}_{N})\|$ is smaller for $z$ closer to 0.
Refer to caption Refer to caption
(a) 2e-14 (b) 2.5e-14
Refer to caption Refer to caption
(c) 1e-13 (d) 1e-12
[…]
(a) dictionary size: 441
(b) dictionary size: 600
(c) dictionary size: 806
Figure 2: Effect of increasing the dictionary size NN (number of basis functions) on the detection of spectral points.
(a) gEDMD
(b) Threshold: 1e-14
(c) Threshold: 1e-13
(d) Threshold: 1e-12
Figure 3: Evaluation of $1/\|\mathbf{R}(x,\mathbf{A}_{N})\|$ at $x=\pm 0.01$.
6.2 Comparing pseudospectra contour
@@ 6.3 Clustering two oscillators with thermal noise
Throughout this example we take $w_{m}=1/M$ for randomly sampled snapshots [klus2015numerical, korda2018convergence].
@@
To compare the dynamic features extracted by ResDMD and Resolvent DMD, we follow the clustering framework in [sakata2024enhancing]. For each method, we collect a family of pseudoeigenfunctions $g_{(z)}$} over a grid of $z$ in the complex plane. We then compute kernel principal angles between the subspaces spanned by these pseudoeigenfunctions, construct a similarity matrix from these angles, and apply spectral clustering to obtain a low-dimensional feature representation. Finally, we perform fuzzy C-means clustering in this embedded feature space to account for ambiguity induced by continuous spectra and obtain three clusters that we interpret as candidate dynamic components.
@@
To compare the dynamic features extracted by ResDMD and Resolvent DMD, we follow the clustering framework in [sakata2024enhancing]. For each method, we collect a family of pseudoeigenfunctions $g_{(z)}$} over a grid of $z$ in the complex plane. We then compute kernel principal angles between the subspaces spanned by these pseudoeigenfunctions, construct a similarity matrix from these angles, and apply spectral clustering to obtain a low-dimensional feature representation. Finally, we perform fuzzy C-means clustering in this embedded feature space to account for ambiguity induced by continuous spectra and obtain three clusters that we interpret as candidate dynamic components.
To connect the clustering in the embedded space with spectral information, we compute spectral measures from the pseudoeigenfunctions using the rigorous resolvent-based procedure developed in [colbrook2024rigorous]. Figure 6 shows the resulting spectral measures colored by cluster labels for ResDMD (left) and Resolvent DMD (right). In the Resolvent DMD case, the two broad peaks associated with the physical oscillation frequencies are each assigned predominantly to a distinct cluster, and a third cluster captures the remaining higher-frequency components. Thus the cluster structure aligns well with the two physical spectral peaks. In the ResDMD case, however, the cluster assignment on the low-frequency side is substantially mixed: the spectral contribution around the lower physical peak is not cleanly separated from nearby components, and the resulting clusters no longer correspond clearly to the two physical modes. This suggests that, when two physical modes occupy nearby spectral regions, Resolvent DMD has higher discriminative power than ResDMD in separating those modes, even though both methods share the goal of mitigating spectral pollution.
@@
Figure 6: Clustering of spectral measures for ResDMD (left) and Resolvent DMD (right), computed using the resolvent-based procedure of [colbrook2024rigorous]. Resolvent DMD separates the two physical peaks into distinct clusters, whereas ResDMD does not clearly isolate the lower-frequency peak.
@@ 7 Conclusion
In this work, we have presented a novel data-driven framework for the spectral analysis of the Koopman operator through the lens of pseudo-resolvents. By leveraging the Sherman-Morrison-Woodbury identity, we constructed a direct matrix approximation of the resolvent operator based on Galerkin approximations for spectral detection.Our theoretical analysis establishes that the proposed method ensures spectral convergence in the Hausdorff metric under the resolvent stability conditions. Furthermore, we derived rigorous error bounds for the arithmetic mean of approximate eigenvalues, controlled by a computable residual quantity $\eta_{N}$. This offers a quantifiable measure of reliability that is often missing in most other heuristic DMD approaches.
@@ Acknowledgment
Y.X. and I.I. acknowledge support from JST CREST Grant Number JPMJCR24Q1 including AIP challenge program, Japan.

### v2

@@ 6.1 Pendulum
In Figure 2, we compare the Koopman operator $\mathbf{K}_{N}$ approximated via standard EDMD against the resolvent operator $\mathbf{R}(z,\mathbf{K}_{N})$ computed using our method. We utilize dictionary sizes of 225, 361, 600, and 702 functions on a separate simulated dataset, testing thresholds of $2.5\times 10^{-14}$ and $3.5\times 10^{-14}$. The probe points $z$ are distributed along $\Gamma_{1.01}$, a slight perturbation of the unit circle chosen to avoid numerical singularities. As the dictionary size $N$ increases, the indicator $1/\|\mathbf{R}(z,\mathbf{K}_{N})\|$ diminishes for a larger proportion of grid points. This leads to a higher detection rate along $\Gamma_{1.01}$ for a fixed threshold, which suggests a clear convergence behavior as the approximation subspace grows.
In Figure 3, we further evaluate the continuous-time Koopman generator using the pseudo-resolvent matrix from Eq. (6). Consistent with the conservative nature of the system described at the outset, we verify whether the detected spectrum aligns with the imaginary axis. To avoid singularities, we sample the resolvent indicator on two vertical lines $\Re(z)=\pm 0.01$ flanking the imaginary axis. Using the criterion $1/\|\mathbf{R}(z,\mathbf{A}_{N})\|<\varepsilon$, we observe that as $\varepsilon$ is relaxed from $10^{-14}$ to $10^{-12}$, the detected set expands but remains tightly concentrated near the imaginary axis. This accurately reflects the physically expected purely oscillatory behavior without introducing artificial damping.
For comparison, Figure 4 displays the generator spectrum obtained via gEDMD [17]. In the gEDMD approach (or equivalently, mapping discrete EDMD eigenvalues $\mu_{j}$ via the principal logarithm $\lambda_{j}\approx\Delta t^{-1}\log\mu_{j}$), the spectrum exhibits a significant spread into the left half-plane. This phenomenon indicates spurious dissipation: the finite-dimensional truncation causes the eigenvalues to contract inside the unit circle ($|\mu_{j}|<1$), which the logarithm translates into negative real parts ($\Re(\lambda_{j})<0$). This artificial decay violates the conservation of energy inherent to the pendulum. In contrast, our pseudo-resolvent approach avoids the direct matrix logarithm and successfully preserves the Hamiltonian spectral structure along the imaginary axis.
Finally, Figure 5 presents a sensitivity analysis regarding the detection threshold $\varepsilon$ over the range $10^{-16},10^{-10}]$. The curve tracks the number of detected spectral points out of a total of 60 test probes located along two vertical lines $\Re(z)=\pm 0.01$, designed to probe the generator spectrum along the imaginary axis. We observe a sharp phase transition in the detection count: below the threshold $10^{-14}$, no points are detected; however, as $\varepsilon$ increases, the count rises rapidly until reaching a saturation status where all 60 points are successfully identified (e.g., $0/60\rightarrow 39/60\rightarrow 60/60$). This saturation indicates that the physical spectrum is well-separated from the underlying numerical errors without the need for precise threshold tuning. Collectively, these results demonstrate that the computed singular values exhibit a clear separation between spectral and non-spectral regions, which confirms that the resolvent norm criterion provides a robust diagnostic for identifying the generator spectrum directly from data.
Figure 1: Koopman spectrum of the nonlinear pendulum: EDMD eigenvalues (black) and resolvent-based detection (red) under successive threshold refinement.
Figure 2: Effect of dictionary size on resolvent-based Koopman spectrum detection for the nonlinear pendulum: larger dictionary size gives higher detection rates under two different singular-value thresholds.
Figure 3: Resolvent-based detection of the Koopman generator spectrum along the imaginary axis under successive threshold refinement ($\epsilon=10^{-14},10^{-13},10^{-12}$).
Figure 4: Spectrum of the EDMD-approximated Koopman generator for the nonlinear pendulum by taking the matrix logarithm of the Koopman operator.
Figure 5: Sensitivity of spectrum detection to the varying singular-value threshold $\epsilon$: a sharp transition from 0 to full detection.
6.2 Linear oscillator
Consider the two-dimensional linear oscillator
$\dot{x}=Ax\quad\text{with}\quad A=\begin{pmatrix}-0.3&-2\\ 2&-0.3\end{pmatrix}$,
discretized at $\Delta t=0.2$. The Koopman eigenvalues of this system are analytically known: $\lambda_{n_{1},n_{2}}=z_{1}^{n_{1}}z_{2}^{n_{2}}$, where $z_{1,2}=e^{(-0.3\pm 2i)\cdot 0.2}$ and $|z_{1}|\approx 0.9418$. We approximate the Koopman operator via EDMD with a Random Fourier Feature (RFF) [28] dictionary of size $N\in\{20,30,50,80,120,180,250,350,500\}$ with kernel length-scale $\ell=1.5$, giving the finite-section operator $\mathcal{K}_{N}=\Psi_{X}^{+}\Psi_{Y}\in\mathbb{R}^{N\times N}$.
Theorem 5.4 asserts that, for each isolated eigenvalue $\lambda$ of the Koopman operator $\mathcal{K}$ with finite algebraic multiplicity $m_{\lambda}$ (i.e., $|\lambda-\bar{\lambda}_{N}|\;\leq\;C\,\eta_{N}$), where $\bar{\lambda}_{N}=m_{\lambda}^{-1}\sum_{j=1}^{m_{\lambda}}\mu_{N,j}$ is the arithmetic mean of the EDMD eigenvalues clustering around $\lambda$ (for a simple eigenvalue, simply $\mu_{N}$), $C$ is a constant independent of $N$, and $\eta_{N}$ is the operator-level truncation error defined in Definition 5.2 (i.e., $\eta_{N}=\|(\mathcal{K}-\mathcal{K}_{N})\,\mathcal{P}_{N}^{\lambda}\|$) with $\mathcal{P}_{N}^{\lambda}$ the spectral projection of $\mathcal{K}_{N}$ associated with the eigenvalues near $\lambda$.
By definition, the Koopman operator acts as $\mathcal{K}f(x)=f(T(x))$ for any observable $f$, where $T$ is the discrete-time flow map. For a general nonlinear system, $T$ is typically unknown and only accessible through data, making direct evaluation of $(\mathcal{K}-\mathcal{K}_{N})$ infeasible. Since the flow map $T(x)=Mx$, $M=e^{A\Delta t}$ of the linear oscillator is known in closed form, we can evaluate $\mathcal{K}f(x)=f(Mx)$ at any point $x$ for any observable $f$. This makes it possible to compute the truncation error $(\mathcal{K}-\mathcal{K}_{N})$ on any dictionary function directly, with no approximation beyond Monte Carlo quadrature.
For each target eigenvalue $\lambda$ (all five are simple, $m_{\lambda}=1$), let $v_{N}$ and $w_{N}$ denote, respectively, the right and left eigenvectors of the EDMD matrix $K_{N}$, normalized so that $w_{N}^{H}v_{N}=1$. The approximate eigenfunction is $g_{N}(x)=\sum_{j}(v_{N})_{j}\,\psi_{j}(x)$, and the truncation error acts as
$(\mathcal{K}-\mathcal{K}_{N})\,g_{N}(x)\;=\;g_{N}(Mx)-\mu_{N}\,g_{N}(x),$
since $\mathcal{K}_{N}g_{N}=\mu_{N}g_{N}$ by construction, whereas $\mathcal{K}g_{N}(x)=g_{N}(Mx)\neq\mu_{N}g_{N}(x)$ (the dictionary functions are not Koopman eigenfunctions). For a rank-one spectral projection $\mathcal{P}_{N}^{\lambda}$, the operator norm evaluates to
$\eta_{N}\;=\;\sqrt{w_{N}^{H}G^{-1}w_{N}}\;\cdot\;\|(\mathcal{K}-\mathcal{K}_{N})\,g_{N}\|_{L^{2}(\mu)},$ (11)
where $G$ is the Gram matrix of the dictionary in $L^{2}(\mu)$. We estimate both the $L^{2}$ norm and $G$ via Monte Carlo integration with $10^{5}$ i.i.d. samples drawn from the same distribution as the training data. Here $G^{-1}$ is computed via a (regularized) pseudoinverse or Tikhonov regularization when needed.
Since $C$ in Eq.(10) is unknown but $N$-independent, the theorem predicts that the ratio $|\mu_{N}-\lambda|/\eta_{N}$ should remain bounded as $N\to\infty$. We therefore track this ratio across all dictionary sizes and modes as the verification strategy. For each dictionary size $N$ and each of the five target modes $\{(1,0),(0,1),(2,0),(0,2),(1,1)\}$, we identify the EDMD eigenvalue $\mu_{N}$ closest to the true eigenvalue $\lambda$ and compute $\eta_{N}$ via Eq.(11).
Figure 6 plots the eigenvalue error $|\mu_{N}-\lambda|$ against $\eta_{N}$ on a log–log scale. All 45 data points ($9$ dictionary sizes $\times$ 55 modes) lie strictly below the diagonal $|\mu_{N}-\lambda|=\eta_{N}$, and also below the fitted line $|\mu_{N}-\lambda|=C\,\eta_{N}$ with $C=0.04$. Both $|\mu_{N}-\lambda|$ and $\eta_{N}$ decrease as $N$ grows (visible from the annotations for mode $(1,0)$), and the two quantities decrease in a correlated fashion, as the theorem predicts.
Figure 7 provides a complementary view. The left panel shows the convergence of both $|\mu_{N}-\lambda|$ (solid) and $\eta_{N}$ (dashed) as functions of $N$ for each mode; more specifically, the error is consistently several orders of magnitude smaller than $\eta_{N}$, and both decline with increasing $N$. The right panel plots the ratio $|\mu_{N}-\lambda|/\eta_{N}$ versus $N$. Across all modes and dictionary sizes, the ratio remains bounded above by $C_{\max}\approx 0.04$, with no upward trend as $N$ increases. The maximum observed ratio is
$\max_{N,\,\lambda}\;\frac{|\mu_{N}-\lambda|}{\eta_{N}}\;=\;0.042\;\ll\;1,$
which confirms the existence of a finite, $N$-independent constant $C$ as asserted by Theorem 5.4.
The fact that $C_{\max}\ll 1$ indicates that the bound is conservative; more specifically, the eigenvalue error is roughly 25 times smaller than the truncation error $\eta_{N}$. This is indeed expected, because the theorem provides a worst-case guarantee over all directions in the spectral subspace, while the actual perturbation $(\mathcal{K}-\mathcal{K}_{N})$ projects only weakly onto the most sensitive direction of the resolvent.
Figure 6: Eigenvalue error $|\mu_{N}-\lambda|$ versus truncation error $\eta_{N}=\|(\mathcal{K}-\mathcal{K}_{N})\mathcal{P}_{N}^{\lambda}\|$ for five Koopman modes across nine dictionary sizes. The dashed diagonal marks $|\mu_{N}-\lambda|=\eta_{N}$; the red line shows the fitted bound with $C=0.04$. All points lie well below the diagonal, which confirms Eq. (10).
Figure 7: Left: convergence of $|\mu_{N}-\lambda|$ (solid) and $\eta_{N}$ (dashed) as $N$ increases. Right: ratio $|\mu_{N}-\lambda|/\eta_{N}$ versus NN; the dashed grey line marks the overall maximum $C_{\max}=0.04$. The ratio remains bounded, consistent with Theorem 5.4.
6.3 Comparison of pseudospectral contours
@@ 6.4 Clustering two oscillators with thermal noise
Throughout this example we take $w_{m}=1/M$ for randomly sampled snapshots [16, 20].
@@
To compare the dynamic features extracted by ResDMD and Resolvent DMD, we follow the clustering framework in [29]. For each method, we collect a family of pseudoeigenfunctions $\{g{(z)}\}$ over a grid of $z$ in the complex plane. We then compute kernel principal angles between the subspaces spanned by these pseudoeigenfunctions, construct a similarity matrix from these angles, and apply spectral clustering to obtain a low-dimensional feature representation. Finally, we perform fuzzy C-means clustering in this embedded feature space to account for ambiguity induced by continuous spectra and obtain three clusters that we interpret as candidate dynamic components.
@@
To connect the clustering in the embedded space with spectral information, we compute spectral measures from the pseudoeigenfunctions using the rigorous resolvent-based procedure developed in [5]. Figure 10 shows the resulting spectral measures colored by cluster labels for ResDMD (left) and Resolvent DMD (right). In the Resolvent DMD case, the two broad peaks associated with the physical oscillation frequencies are each assigned predominantly to a distinct cluster, and a third cluster captures the remaining higher-frequency components. Thus the cluster structure aligns well with the two physical spectral peaks. In the ResDMD case, however, the cluster assignment on the low-frequency side is substantially mixed: the spectral contribution around the lower physical peak is not cleanly separated from nearby components, and the resulting clusters no longer correspond clearly to the two physical modes. This suggests that, when two physical modes occupy nearby spectral regions, Resolvent DMD has higher discriminative power than ResDMD in separating those modes, even though both methods share the goal of mitigating spectral pollution.
@@
Figure 10: Clustering of spectral measures for ResDMD (left) and Resolvent DMD (right), computed using the resolvent-based procedure of [5]. Resolvent DMD separates the two physical peaks into distinct clusters, whereas ResDMD does not clearly isolate the lower-frequency peak.
@@ 7 Conclusion
In this work, we have presented a novel data-driven framework for the spectral analysis of the Koopman operator through the lens of pseudo-resolvents. By leveraging the Sherman-Morrison-Woodbury identity, we constructed a direct matrix approximation of the resolvent operator based on Galerkin approximations for spectral detection. Our theoretical analysis establishes that the proposed method ensures spectral convergence in the Hausdorff metric under the resolvent stability conditions. Furthermore, we derived rigorous error bounds for the arithmetic mean of approximate eigenvalues, controlled by a residual quantity $\eta_{N}$. This offers a quantifiable measure of reliability that is often missing in most other heuristic DMD approaches.
@@ Acknowledgment
Y.X. and I.I. acknowledge support from JST CREST Grant Number JPMJCR24Q1 including AIP challenge program, Japan. I.S. acknowledges support from JSPS KAKENHI Grant Number JP24K20864.